Find the three consecutive numbers such that the sum of double the first and the third totals 362
Found 2 solutions by checkley75, cheesey:
Answer by checkley75(3666)

x, x+1, x+2 are the three consecutive numbers.
2x+(x+2)=362
2x+x+2=362
3x=362-2
3x=360
x=360/3
x=120 answer for the first number.
x+2=122 answer for the third number.
proof:
2*120+122=362
240+122=362
362=362

Answer by cheesey(22)

x, x+1, x+2
2x + x+2 = 362
3x + 2 = 362
3x = 360
x=120
thus (120, 121, 122)